solve 4x^2 + 12x - 7=0 by completing the square.
Answer by stanbon(75887)

solve 4x^2 + 12x - 7=0 by completing the square
----------------
4(x^2 + 3x + ?) = 7 + 4*?
---
4(x^2 + 3x + (3/2)^2) = 7 + 4(3/2)^2
---
4(x+(3/2))^2 = 16
---
(x+(3/2)) = +-sqrt(4)
(x+(3/2) = +-2
-----
x = 1/2 or x = -7/2
--------------------------
Cheers,
Stan H.
Reply to stanbon@comcast.net